OpenStudy (anonymous):

what is the equation (not the answer) for the total arc length of an ellipse when x=4sin(theta) and y=3cos(theta)

OpenStudy (anonymous):

line integral integrate from 0 to 2 pi ;Sqrt[(x')^2 +(y')^2]

OpenStudy (anonymous):

Its multiple choice and there are two answers with that, one where the whole integral is multplied by for and the other is multiplied by 16

OpenStudy (anonymous):

4* not for

OpenStudy (anonymous):

oh, that just symmetry

OpenStudy (anonymous):

what are the limits?

OpenStudy (anonymous):

they are both 0 to 2pi

OpenStudy (anonymous):

what was the integrand

OpenStudy (anonymous):

Sqrt[cos^2(x)+9/16 Sin^2(x)] ?

OpenStudy (anonymous):

\[16\int\limits_{0}^{2\pi}\sqrt{16\sin^{2}\theta+9\cos^{2}\theta} d \theta\] or \[4\int\limits_{0}^{2\pi}\sqrt{16\sin^{2}\theta+9\cos^{2}\theta} d \theta\]

OpenStudy (anonymous):

I don't know where 4 and 16 are coming from because we are not using symetry

OpenStudy (anonymous):

thats what im confused about too.

OpenStudy (anonymous):

there are 4 answers. all of them have the same thing under the radical, its the limits and the number its multiplied by that are varied

OpenStudy (anonymous):

put all of them up

OpenStudy (anonymous):

others are o to theta...

OpenStudy (anonymous):

a) \[4\int\limits_{0}^{\pi/2}\] b) \[16\int\limits_{0}^{2 \pi}\] c)\[16\int\limits_{0}^{ \pi}\] d)\[4\int\limits_{0}^{2 \pi}\]

OpenStudy (anonymous):

a) because we are doing quater and using symmtary

OpenStudy (anonymous):

thats what i thought too, but that was wrong

OpenStudy (anonymous):

are you absolutely sure the integrand are same, lol because other than that I think we are right

OpenStudy (anonymous):

yes i am positive haha.

OpenStudy (anonymous):

I have had 3 different people look at it and we all come to the same conclusion so im thinking maybe its a mistake on the problem

OpenStudy (anonymous):

I used to see it as a sqrt (1- sin^2t) 0 to theta

OpenStudy (anonymous):

The 1 - sin^2 is just a trig rewrite...

OpenStudy (anonymous):

k sin^2, sorry...

OpenStudy (anonymous):

estudier that is only if y is function of x

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x=4sin(theta) and y=3cos(theta), right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

\[\int \sqrt{1+(dy/dx)^2}\] ?

OpenStudy (anonymous):

No, I mean u can use sin^2 + cos^2 = 1 to rewrite....

OpenStudy (anonymous):

there should be no number in front of integral

OpenStudy (anonymous):

Anyway, we are talking about the same thing just restricted to certain angles..

OpenStudy (anonymous):

there are two links up the , btw

OpenStudy (anonymous):

They have a 12 and a 16 in the answers.

OpenStudy (anonymous):

I meant the numerical answer 22 is same thing for letting wolfram figure out arc length based on equation and letting it figure out integral

OpenStudy (anonymous):

Agree but E is the elliptic integral,,,

OpenStudy (anonymous):

Which is what u are looking for here, right?

OpenStudy (anonymous):

looking for arclength of ellipse,

OpenStudy (anonymous):

That is the elliptic integral....

OpenStudy (anonymous):

Based on the Wolfram links, I would go with the 16...

OpenStudy (anonymous):

but we get same result for \[\int\limits_{0}^{2\pi}\sqrt{16\sin^{2}\theta+9\cos^{2}\theta} d \theta\]=22.1 letting wolfram figure out arclength based on ellipse equation =22.1

OpenStudy (anonymous):

It's the diff between E(7/16) and E(-7/9) which must be to do with the limits...

OpenStudy (anonymous):

See, if u rewrite the cos^2 + sin^2 to sin^2 only, u get the 7.

OpenStudy (anonymous):

but numerical answer is same which must mean the \[\int\limits_{0}^{2\pi}\sqrt{16\sin^{2}\theta+9\cos^{2}\theta} d \theta\] is correct

OpenStudy (anonymous):

Yes and the answer is 16* elliptic integral so u need the 16.

OpenStudy (anonymous):

oh, he is doing elliptical integral?

OpenStudy (anonymous):

The arc length is the elliptic integral...

OpenStudy (anonymous):

The 12 E(-7/9) or 16E(7/16) gives same answer....

OpenStudy (anonymous):

He is just trying to set up a line integral on ellipse

OpenStudy (anonymous):

sqrt(16 sin^2 +9 cos^2) sqrt(7 sin^2 +9) -> 3* int ]sqrt(1+ 7/9 sin^2)] is the elliptic integral (or arc length integral if u prefer)

OpenStudy (anonymous):

Wait a minute, have I got sin and cos backwards?

OpenStudy (anonymous):

If so, maybe it will come out to 4....

OpenStudy (anonymous):

It does come out to 4 if sin and cos are reversed....

OpenStudy (anonymous):

Is that what the problem is..?

OpenStudy (anonymous):

Back to the beginning, let me look....

OpenStudy (anonymous):

Isn't the usual parameters for ellipse (a cos t, b sin t)?

OpenStudy (anonymous):

yes, but it works too

OpenStudy (anonymous):

Yes but the answers in the options are based on the standard paramaterization..

OpenStudy (anonymous):

Yes, That's what I was trying to say

OpenStudy (anonymous):

If u reverse them, then everything will work and the answer is 4*

OpenStudy (anonymous):

So the question is wrong...

OpenStudy (anonymous):

Not wrong, inconsistent with the proposed answers...

OpenStudy (anonymous):

It should be x=4cos(theta) and y=3sin(theta) to match the answers.

OpenStudy (anonymous):

imranmeah, u agree?

OpenStudy (anonymous):

well ive been discussing it with my friend and we have also come to the conclusion that the problem is wrong

OpenStudy (anonymous):

For parametric curve we want the integral of sqrt(x'^2 + y'^2) and if this is sqrt(16 sin^2 + 9 cos^2) then x' = 4 sin^2 -> x = 4 cos and y = 3 sin

OpenStudy (anonymous):

4 sin, sorry, not 4 sin^2

OpenStudy (anonymous):

For paramteric curve we want the integral of sqrt(x'^2 + y'^2) and if this is sqrt(16 sin^2 + 9 cos^2) then x' = 4 sin -> x = 4 cos and y = 3 sin